OpenStudy (anonymous):

What is the sum of a 6-term geometric sequence if the first term is 11, the last term is –11,264 and the common ratio is –4?

OpenStudy (anonymous):

What formula do i use?

OpenStudy (anonymous):

use Sn = a * ( 1 - r^n) ---------- 1 - r

OpenStudy (anonymous):

a - first term r = common ratio, n = no. of terms you dont need to last trm

OpenStudy (anonymous):

* you dont need to use the last term

OpenStudy (anonymous):

Sn = 11 * ( 1 - (-4)^6) / 1-(-4)

OpenStudy (anonymous):

ok ?

OpenStudy (anonymous):

Let me write and solve it on my paper

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

34,375

OpenStudy (anonymous):

oh i got -9009

OpenStudy (anonymous):

(-4)^6 = 4096 right?

OpenStudy (anonymous):

Sn=11*(1-(-4)^6/(1-(-4) Sn= 11* 5^6 / 5 Sn= 11* 15625 /5 Sn= 11* 3125 Sn= 34,375 thats how i got it maybe i did something wrong

OpenStudy (anonymous):

yes - the second line is wrong 1 - (-4)^6 = 1 - 4096 only the (-4 )is taken to the power 6

OpenStudy (anonymous):

note where the parentheses are

OpenStudy (anonymous):

Okay i think i got it

OpenStudy (anonymous):

in the formula the ^n refers to the r only

OpenStudy (anonymous):

good

OpenStudy (anonymous):

how do i solve this one What is the sum of an 8-term geometric sequence if the first term is 10 and the last term is 781,250?

OpenStudy (anonymous):

ok i havent done one of these for a while.... the nth term of a geometric sequence = a r^(n-1) so the eighth term = 10* r^(8-1) 10*r^7 = 781,250 r^7 = 78125 r = 5 so now we know a = 10, r = 5 and n = 8 and we can find the sum of 8 term using the formula S8 = a * r^n - 1 ------- r - 1 notice that r^n comes before the 1 this time because r is positive

OpenStudy (anonymous):

11607.?